Here is the Crew 189 Sol Summary Report 05 March 2018
Summary Title: Beginning of last martian week
Author: Victoria Da-Poian
Mission Status: All nominal
Sol Activity Summary: We began our last week on Mars. Today’s EVA was dedicated to air sensors experiments and to localization experiment. The EVA was not really successful (cf EVA report) but we still have one week left to work on this! In the Hab today, I was pleased to cook a salad (salad, cucumbers, cherry tomatoes!): it was delicious! I took care of the greenhab as Jérémy was outside and as the previous Crew 175 Greenhab officer, I spent some time enjoying the vegetables smells! Then, we spent time working on our experiments (human factors experiment, virtual reality filming…)
Look Ahead Plan: Tomorrow, we will welcome two French television journalists (TF1). They will spend the day with us from our physical training to our afternoon activities.
Anomalies in work: None
Crew Physical Status: Fine
Reports to be filed:
Support Requested: None
You must be logged in to post a comment.